High West, CPAs currently provide tax return preparation services to individuals in the Texas area. Jim...

  1. High West, CPAs currently provide tax return preparation services to individuals in the Texas area. Jim Bean have received a one-time only request to prepare 100 tax returns for clients of another CPA firm in Texas while the owner recovers from major surgery. Jim Bean have the capacity to prepare 700 returns. Jim Bean have an effective tax rate of 30%. Jim Bean income statement before consideration of the one-time request is as follows:

                                    

Sales (500 returns)

$150,000

Variable preparation costs

20,000

Contribution margin

130,000

Fixed preparation costs

80,000

Operating income

50,000

Income taxes

15,000

Net income

$35,000

Required:

In negotiating a price for the special request, what is the minimum amount Jim Bean would accept per return?
Answer #1

Since spare capacity exists, no additional fixed costs will be incurred in the preparation of returns

Minimum price is equal to the variable cost incurred

= 20,000/500

= $40 per return

For whole one tie request of 100 Minimum price = 40*100 = $400